OpenStudy (anonymous):

how do I solve a pythagorean theorem problem when one of the sides is the square root of 15?

OpenStudy (anonymous):

The same way you'd solve it if one of the sides was 4. Just treat the \(\sqrt{15}\) like any other number. (Cause it is)

OpenStudy (anonymous):

Won't be a problem once u have squared it...

OpenStudy (anonymous):

once i square it I get 15

OpenStudy (anonymous):

There u go then...

OpenStudy (anonymous):

ok, thanks! i was making it harder than it is

OpenStudy (anonymous):

what if the number is 3 square root of 2?

OpenStudy (anonymous):

\[(3\sqrt{2})^2 = (3\cdot \sqrt{2})^2 = 3^2 \cdot \sqrt{2}^2 = 9 \cdot 2 = 18\]

OpenStudy (anonymous):

great

OpenStudy (anonymous):

when you have 3x squared times y, do you square the x, then multiply by 3, then multiply by y? Or multiply 3 by x then square?